OpenStudy (anonymous):

sin((17π)/(12)) find the exact value! please!

OpenStudy (anonymous):

i think it is sqrt 6 - sqrt 2 / 4

jimthompson5910 (jim_thompson5910):

do you have a calculator?

OpenStudy (anonymous):

no...

jimthompson5910 (jim_thompson5910):

that's ok, you can use google as a calculator so if you type "sin(17pi/12)" without quotes into google, you would get this http://www.google.com/search?hl=&q=sin%2817pi%2F12%29&sourceid=navclient-ff&rlz=1B3GGLL_enUS420US420&ie=UTF-8

OpenStudy (anonymous):

ok?

jimthompson5910 (jim_thompson5910):

now you type your supposed answer "(sqrt(6)-sqrt(2))/4" without quotes into google to get http://www.google.com/search?hl=&q=%28sqrt%286%29-sqrt%282%29%29%2F4&sourceid=navclient-ff&rlz=1B3GGLL_enUS420US420&ie=UTF-8 Now IF the two answers are the same, then the two expressions are equal to each other

jimthompson5910 (jim_thompson5910):

So let's say hypothetically that sin(17pi/12) gives a result of 0.5 (it doesn't, but let's say it does) IF (sqrt(6)-sqrt(2))/4 also gave a result of 0.5 (it doesn't, but again let's say it does), then we could say that sin(17pi/12) = (sqrt(6)-sqrt(2))/4 so this is how you can check any possible answers you get

terenzreignz (terenzreignz):

Or you could use identities? :D

jimthompson5910 (jim_thompson5910):

so according to the first link I gave you, what is sin(17pi/12)

jimthompson5910 (jim_thompson5910):

true, that's probably the better way to do it

OpenStudy (anonymous):

-0.9659

jimthompson5910 (jim_thompson5910):

now look at the second link, and tell me what the value of (sqrt(6)-sqrt(2))/4 is

terenzreignz (terenzreignz):

\[\Large \frac{17\pi}{12}= \frac{7\pi}{6}+ \frac{\pi}4\] (not related to @jim_thompson5910 's checking method, strictly speaking, but may be useful later on)

OpenStudy (anonymous):

it didnt work!

jimthompson5910 (jim_thompson5910):

look at the second link, what is the value of (sqrt(6)-sqrt(2))/4

OpenStudy (anonymous):

you didnt post a second link!

jimthompson5910 (jim_thompson5910):

yes I did, but it's ok, I'll do it again http://www.google.com/search?hl=&q=%28sqrt%286%29-sqrt%282%29%29%2F4&sourceid=navclient-ff&rlz=1B3GGLL_enUS420US420&ie=UTF-8

OpenStudy (anonymous):

no they didnt match.

jimthompson5910 (jim_thompson5910):

so that proves that sin(17pi/12) = (sqrt(6)-sqrt(2))/4 is false

jimthompson5910 (jim_thompson5910):

so use the identity terenzreignz posted and use other trig identities to simplify and try again

OpenStudy (anonymous):

ok sqrt 6 +sqrt 2 / 4?

jimthompson5910 (jim_thompson5910):

type that into google and tell me what you get

terenzreignz (terenzreignz):

/smh/ If you still care for it, here's the identity you need... \[\Large \sin(\color{red}\alpha+\color{blue}\beta) = \sin(\color{red}\alpha)\cos(\color{blue}\beta) + \cos(\color{red}\alpha) \sin (\color{blue}\beta)\]

OpenStudy (anonymous):

i think 2.80 can you confirm that?

jimthompson5910 (jim_thompson5910):

that's incorrect, you typed in (sqrt(6)+sqrt(2))/4 right?

jimthompson5910 (jim_thompson5910):

and yes, you should use the latest identity terenzreignz posted from there, use the unit circle to evaluate

OpenStudy (anonymous):

the / is divide right?

jimthompson5910 (jim_thompson5910):

yes

jimthompson5910 (jim_thompson5910):

1/2 means 1 divided by 2

OpenStudy (anonymous):

none of my answers match..

jimthompson5910 (jim_thompson5910):

btw, you have to enclose sqrt(6)+sqrt(2) in a set of parenthesis to make sure you divide all of sqrt(6)+sqrt(2) by 4

jimthompson5910 (jim_thompson5910):

sqrt(6)+sqrt(2)/4 is NOT the same as (sqrt(6)+sqrt(2))/4

OpenStudy (anonymous):

this is helping im more lost now then i was before!

OpenStudy (anonymous):

is is -sqrt6 - sqrt2 /4

terenzreignz (terenzreignz):

Guessing when given a multiple choice question may be tempting, but it is ill-advised. Please try to use the identities ~

OpenStudy (anonymous):

look i dont understand the identites you havnt helped me at all!

jimthompson5910 (jim_thompson5910):

here's a more simpler example if I said 1/2 = 2/4 were true, then we would have to show that 1/2 and 2/4 equal the same number since 1/2 turns into 0.5 and 2/4 turns into 0.5 as well (use a calculator), this means that 1/2 = 0.5 2/4 = 0.5 this proves that 1/2 = 2/4 is true

jimthompson5910 (jim_thompson5910):

see how this logic works out?

OpenStudy (anonymous):

yes but i put all my quetions into the calculator and none of them came up! this is timed too!

terenzreignz (terenzreignz):

You want the answer?

jimthompson5910 (jim_thompson5910):

oh so a test?

jimthompson5910 (jim_thompson5910):

sry can't help you there

OpenStudy (anonymous):

no a review.

OpenStudy (anonymous):

but ive had the page open too long.

terenzreignz (terenzreignz):

Well, you answered MY question correctly, at least. Use the identity. It was posted, just a few posts ago.... and take \[\large \color{red}\alpha = \frac{7\pi}{6}\] and \[\large \color{blue}\beta = \frac{\pi}{4}\]

jimthompson5910 (jim_thompson5910):

ok I'm glad the concept of what I'm explaining makes sense that's a start the thing that's probably going wrong is that you're probably entering the expressions incorrectly you are NOT typing this in sqrt(6)+sqrt(2)/4 since you SHOULD be typing this in (sqrt(6)+sqrt(2))/4 the difference is that set of parenthesis that changes everything

OpenStudy (anonymous):

i did that!

jimthompson5910 (jim_thompson5910):

(sqrt(6)+sqrt(2))/4 says to divide all of sqrt(6)+sqrt(2) by 4 sqrt(6)+sqrt(2)/4 says to just divide sqrt(2) by 4

jimthompson5910 (jim_thompson5910):

hmm what could be going wrong then...

OpenStudy (anonymous):

ok i really think its sqrt 2 - sqrt 6 / 4

jimthompson5910 (jim_thompson5910):

oh and sqrt 2 should be sqrt(2) same for sqrt 6

jimthompson5910 (jim_thompson5910):

not sure why the parenthesis aren't making it

OpenStudy (anonymous):

can you check it since you seem to know how to use it!

jimthompson5910 (jim_thompson5910):

but check that with a calculator and tell me what you get

jimthompson5910 (jim_thompson5910):

well I want you show you how to use it so you can check other problems on your own when you get a possible solution, you need to know how to check it with a caclulator

jimthompson5910 (jim_thompson5910):

calculator*

OpenStudy (anonymous):

that thing doenst work ill never useit again .

OpenStudy (anonymous):

can you just tell me if thats right please

terenzreignz (terenzreignz):

You have to use something, though. I still recommend using identities, the old fashioned way of doing things.

OpenStudy (anonymous):

@terenzreignz GO AWAY!

jimthompson5910 (jim_thompson5910):

it works, I just wish I knew what you were typing in (exactly) so I can pinpoint where things are going wrong

OpenStudy (anonymous):

ugh your hellpless both of you

terenzreignz (terenzreignz):

You need to be a bit more diplomatic than that.

jimthompson5910 (jim_thompson5910):

fine, be rude